OpenStudy (anonymous):

Given: F(x) = 5x - 6 and G(x) = x - 4

OpenStudy (anonymous):

i need help

OpenStudy (gorv):

what we need to do?????????

OpenStudy (anonymous):

find f*g inverse

OpenStudy (gorv):

f(g(x)) inverse u mean????

OpenStudy (anonymous):

yes that

OpenStudy (gorv):

or just we need to multiply them

OpenStudy (anonymous):

just multiply them

OpenStudy (anonymous):

Given: F(x) = 3x 2 + 1, G(x) = 2x - 3, H(x) = x

OpenStudy (gorv):

(5x-6)(x-6)=5x^2-6x-30x+36=5x^2-36x+36

OpenStudy (gorv):

let y=5x^2-36x+36

OpenStudy (gorv):

now before doing inverse

OpenStudy (gorv):

f(x)=y=x^2+1 how we find inverse

OpenStudy (gorv):

just find value of x y-1=x^2 (y-1)^1/2=x now for inverse just switch x and y

OpenStudy (gorv):

y=(x-1)^1/2 is your inverse

OpenStudy (gorv):

do same for our prob

OpenStudy (gorv):

find value of x y=5x^2-36x+36 5x^2-36x+36=y 5x^2-36x+36-y=0 using quadratic formula\[\frac{ 36\pm \sqrt{36^2-4*5*(36-y)}}{ 10 }\]

OpenStudy (gorv):

that is the value of x

OpenStudy (gorv):

\[x=\frac{ 36\pm \sqrt{576+20y} }{ 10}\]

OpenStudy (gorv):

on switching

OpenStudy (gorv):

\[y=\frac{ 36\pm \sqrt{576+20x} }{ 10 }\]

OpenStudy (gorv):

it is our inverse

OpenStudy (gorv):

similarly do for other

OpenStudy (gorv):

@quashia01